OpenStudy (anonymous):

can someone help me with this problem? 2(w-6.2)+5.4=13.1

OpenStudy (anonymous):

solve for w?

OpenStudy (anonymous):

use order of operations backwards

OpenStudy (anonymous):

to get everything away from the w The first step is to subtract 5.4 from both sides.

OpenStudy (anonymous):

Do you understand how to continue from there?

OpenStudy (anonymous):

okay i think so so like this 15.4=13.1 15.4 15.4

OpenStudy (anonymous):

When you subtract 5.4 from both sides you should get: 2(w-6.2)=7.7

OpenStudy (anonymous):

I don't understand what you did.

OpenStudy (anonymous):

i need like a step by step how to do it. i dont understand how to do this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You start with: 2(w-6.2)+5.4=13.1 You subtract 5.4 from both sides and end up with: 2(w-6.2)=7.7 You divide both sides by 2 and get: w-6.2=3.85 You add 6.2 to both sides to get your final answer of: w=10.05

OpenStudy (anonymous):

ok but show me how u did it.

OpenStudy (anonymous):

order of operations: parentheses exponents multiplication and division addition and subtraction

OpenStudy (anonymous):

I did it backwards because if you start with whatever the value for w is you would first do what's in the parentheses, then multiply then add

OpenStudy (anonymous):

so in order to get w, you have to do it backwards.